# Time-dependent dynamics in the confined lattice Lorentz gas

### v1

@@ Acknowledgments
AS gratefully acknowledges FWF Der Wissenschaftsfonds for funding through the Lise-Meitner Fellowship (Grant DOI 10.55776/M3300). TF gratefully acknowledges support by the Austrian Science Fund (FWF) 10.55776/I5257. TF gratefully acknowledges hospitality of Université Pierre et Marie Curie where parts of the project have been performed.
@@ A.1 Green’s functions on the cylinder
[…]
@@
×$\times\int_{-\pi}^{\pi}\frac{\mathrm{d}k_{x}}{2\pi}e^{\mathsf{i}k_{x}x}\exp[\epsilon_{x}(k_{x})t]\,.$
@@ A.2 Green’s functions with biasing force
where $I_{x}$ is a modified Bessel function of the first kind of order $x$. Taking $\alpha=(t/2)\cosh(F/2)$ and $\beta=(t/2)\sinh(F/2)$, the above formula yields
@@
×$\times\int_{0}^{\infty}\textrm{d}t\,\textrm{e}^{-r_{q}t/2}I_{x}(t/2)\,,$
@@ A.3 Low-frequency expansion
Notice that the first two terms in the above expression correctly recover the smooth part of the corresponding result obtained for the unbounded model [Eq. (V.1)]. In order to recover also the singular term $-(4/\pi)s\ln s$ one can perform a scaling limit such that $s\rightarrow 0$, $L\rightarrow\infty$ with the variable $\tilde{s}=sL^{2}$ fixed; for a detailed discussion on this aspect we refer the reader to Appendix C.4.
@@ C.1 Orthogonal transformation
The transformed basis and the basis $\{\textbf{e}_{-2},\dots,\textbf{e}_{2}\}$ are related to each other via
@@ C.2 Moments of the displacement
where $\langle p_{x}|\hat{t}|n\rangle=t_{25}^{\prime}$. This can be extracted from $t_{25}^{\prime}=(S^{\prime}v^{\prime})_{25}$, namely from the product of the second row of $S^{\prime}$ with the fifth column of $v^{\prime}$, the latter contains only one non-vanishing entry, $v_{25}^{\prime}=2v_{0}/\sqrt{10}$, thus $t_{25}^{\prime}=(2v_{0}/\sqrt{10})S_{25}^{\prime}$. Therefore the function $\mathcal{V}_{L}(F;s)$ defined by $\mathsf{i}\partial_{k_{x}}Nt|_{\textbf{k}=0}:=v_{0}\mathcal{V}_{L}(F;s)$ is given by $\mathcal{V}_{L}(F;s)=-2v_{0}S_{25}^{\prime}$. Although the full dependence on $F$ turns out to be rather long to be reported here, for small $F$ we have
@@
where $\Delta_{L}(s)$ is defined in the main text and in Eq. (V.1) it is evaluated for some values of $L$. For the non confined model ($L\rightarrow\infty$) the result for $\mathcal{V}_{L}(F;0)$ is
@@ C.4 Time-dependent diffusion coefficient
The plot of Fig. 7 provides the response to the time-dependent diffusion coefficient as a function of the rescaled time $t/L^{2}$. The response is measured in units of the asymptotic tail $\Delta D_{\rm tail}(t_{L})$. By inspection of this plot we conclude that for $L\rightarrow\infty$ a scaling function $\mathcal{D}(\tilde{t})$ characterized by two distinct power laws is actually emerging.
@@
$\mathcal{D}(\tilde{t})=\lim_{L\rightarrow\infty}\frac{2L^{2}}{\pi}\int_{0}^{\infty}\textrm{d}\omega\,\textrm{Re}\left(F_{L}(\mathsf{i}\omega)\right)\cos(\omega L^{2}\tilde{t})\,,$ (156)
a representation that follows from Eq. (150) and the scaling ansatz [Eq. (154)]. This integral representation suggests to perform the change of variables $\Omega=\omega L^{2}$ in Eq. (156), the latter yields a rather suggestive form for the scaling function $\mathcal{D}(\tilde{t})$:
$\mathcal{D}(\tilde{t})=\lim_{L\rightarrow\infty}\frac{2}{\pi}\int_{0}^{\infty}\textrm{d}\Omega\,\textrm{Re}\left(F_{L}(\mathsf{i}\Omega/L^{2})\right)\cos(\Omega\tilde{t})\,;$ (157)
however, this representation presents some subtleties due to the simplification of the overall $L^{2}$ outside the integral with the $L^{-2}$ stemming from $\textrm{d}\omega=L^{-2}\textrm{d}\Omega$. This is certainly correct provided $L$ is finite, while this operation must be handled with great care if $L\rightarrow\infty$, in particular also the order of limit and integration must be taken with caution.

### v2

@@ Acknowledgments
This research was funded in part by the Austrian Science Fund (FWF) through the Lise-Meitner Fellowship (Grant DOI 10.55776/M3300) and 10.55776/P35673. TF gratefully acknowledges hospitality of Université Pierre et Marie Curie where parts of the project have been performed.
The following is a series of appendices dedicated to technical details. In particular, Appendix A deals with Green’s function on the cylinder both with and without external force, as well as the low-frequency expansion of propagators. In Appendix B we show how to express the variance of the displacement in terms of quantities accessible in the scattering formalism. Finally, Appendix C provides numerous details involved in the calculations within the scattering formalism.
@@ A.1 Green’s functions on the cylinder
The explicit form of the propagator in the basis of lattice sites $|\textbf{r}\rangle$ can be obtained from the conditional probability in the plane wave basis. The probability to observe the tracer at site r at time $t$ knowing that it started its motion at position $\textbf{r}^{\prime}$ time $t=0$ is
@@
$\times\int_{-\pi}^{\pi}\frac{\mathrm{d}k_{x}}{2\pi}e^{\mathsf{i}k_{x}x}\exp[\epsilon_{x}(k_{x})t]\,.$
@@ A.2 Green’s functions with biasing force
where $I_{x}(\cdot)$ is a modified Bessel function of the first kind of order $x$. Taking $\alpha=(t/2)\cosh(F/2)$ and $\beta=(t/2)\sinh(F/2)$, the above formula yields
@@
$\times\int_{0}^{\infty}\textrm{d}t\,\textrm{e}^{-r_{q}t/2}I_{x}(t/2)\,$,
@@ A.3 Low-frequency expansion
Notice that the first two terms in the above expression correctly recover the smooth part of the corresponding result obtained for the unbounded model [Eq. (V.1)]. In order to recover also the singular term $(4/\pi)s\ln s$ one can perform a scaling limit such that $s\rightarrow 0$, $L\rightarrow\infty$ with the variable $\tilde{s}=sL^{2}$ fixed; for a detailed discussion on this aspect we refer the reader to Appendix C.4.
@@ C.1 Orthogonal transformation
The transformed basis and the basis $(\textbf{e}_{-2},\dots,\textbf{e}_{2})$ are related to each other via
@@ C.2 Moments of the displacement
where $\langle p_{x}|\hat{t}|n\rangle=t_{25}^{\prime}$. This can be extracted from $t_{25}^{\prime}=(S^{\prime}v^{\prime})_{25}$, namely from the product of the second row of $S^{\prime}$ with the fifth column of $v^{\prime}$, the latter contains only one non-vanishing entry, $v_{25}^{\prime}=2v_{0}/\sqrt{10}$, thus $t_{25}^{\prime}=(2v_{0}/\sqrt{10})S_{25}^{\prime}$. Therefore the function $\mathcal{V}_{L}(F;s)$ defined by $\mathsf{i}\partial_{k_{x}}Nt|_{\textbf{k}=0}:=v_{0}\mathcal{V}_{L}(F;s)$ is given by $\mathcal{V}_{L}(F;s)=-2v_{0}S_{25}^{\prime}$. Although the full dependence on $F$ turns out to be rather long to be reported here, provided $s>0$, then for small $F$ we have
@@
where $\Delta_{L}(s)$ is defined in Eq. (80) and in Eq. (V.1) it is evaluated for some values of $L$. For the sake of completeness, we mention that the small-$F$ expansion of $\mathcal{V}_{L}(F;s)$ at $s=0$ contains, beyond the $F$-independent term, a term of order $O(F)$. For the non confined model ($L\rightarrow\infty$) the result for $\mathcal{V}_{L}(F;0)$ is
@@ C.4 Time-dependent diffusion coefficient
[…]
@@
[…]
however, this representation presents some subtleties due to the simplification of the overall $L^{2}$ outside the integral with the $L^{-2}$ stemming from $\textrm{d}\omega=L^{-2}\textrm{d}\Omega$. This is certainly correct provided $L$ is finite, while this operation must be handled with care if $L\rightarrow\infty$, in particular also the order of limit and integration is important.
Appendix D The function $\mathcal{V}_{L}(F;s)$ for the quasi periodic two-lane model ($L=2$)
Here we provide the explicit expression of $\mathcal{V}_{2}(F;s)$. Since the result is rather long, we will write
$\mathcal{V}_{2}(F;s)=\frac{A(\cosh(F),\cosh(F/2),s)}{B(\cosh(F),\cosh(F/2),s)}\,$, (159)
where $A(a,b,s)$ and $B(a,b,s)$ are the functions given by
$A(a,b,s)$ $=\sqrt{2}b^{2}\sqrt{(b+2s-1)(b+2s+3)}\sqrt{a+8s(b+s)-1}+2\sqrt{2}bs\sqrt{(b+2s-1)(b+2s+3)}\sqrt{a+8s(b+s)-1}$
$+4\sqrt{2}b\sqrt{(b+2s-1)(b+2s+3)}\sqrt{a+8s(b+s)-1}-\sqrt{2}\sqrt{(b+2s-1)(b+2s+3)}\sqrt{a+8s(b+s)-1}$
$-4b^{4}-24b^{3}s-12b^{3}-48b^{2}s^{2}-48b^{2}s-4b^{2}\sqrt{(b+2s-1)(b+2s+3)}$
$+2b^{2}\sqrt{(b+2s-1)(b+2s+3)}\sqrt{(b+2s+1)(b+2s+3)}+4b^{2}-32bs^{3}-48bs^{2}+8bs$
$-8bs\sqrt{(b+2s-1)(b+2s+3)}+4bs\sqrt{(b+2s-1)(b+2s+3)}\sqrt{(b+2s+1)(b+2s+3)}$
$+4\sqrt{(b+2s-1)(b+2s+3)}-2\sqrt{(b+2s-1)(b+2s+3)}\sqrt{(b+2s+1)(b+2s+3)}+12b\,$, (160)
and
$B(a,b,s)$ $=2\sqrt{2}b^{2}s\sqrt{a+8s(b+s)-1}+8\sqrt{2}s^{3}\sqrt{a+8s(b+s)-1}+8\sqrt{2}bs^{2}\sqrt{a+8s(b+s)-1}$
$+16\sqrt{2}s^{2}\sqrt{a+8s(b+s)-1}+8\sqrt{2}bs\sqrt{a+8s(b+s)-1}+6\sqrt{2}s\sqrt{a+8s(b+s)-1}$
$-6b^{3}s-4b^{3}-36b^{2}s^{2}-42b^{2}s-2b^{2}s\sqrt{(b+2s-1)(b+2s+3)}+4b^{2}s\sqrt{(b+2s+1)(b+2s+3)}$
$+4b^{2}\sqrt{(b+2s+1)(b+2s+3)}-12b^{2}-72bs^{3}-8s^{3}\sqrt{(b+2s-1)(b+2s+3)}$
$+16s^{3}\sqrt{(b+2s+1)(b+2s+3)}-120bs^{2}-8bs^{2}\sqrt{(b+2s-1)(b+2s+3)}-24s^{2}\sqrt{(b+2s-1)(b+2s+3)}$
$+16bs^{2}\sqrt{(b+2s+1)(b+2s+3)}+32s^{2}\sqrt{(b+2s+1)(b+2s+3)}-42bs-12bs\sqrt{(b+2s-1)(b+2s+3)}$
$-6s\sqrt{(b+2s-1)(b+2s+3)}+24bs\sqrt{(b+2s+1)(b+2s+3)}-4s\sqrt{(b+2s+1)(b+2s+3)}$
$+4b\sqrt{(b+2s+1)(b+2s+3)}-8\sqrt{(b+2s+1)(b+2s+3)}+4b-48s^{4}-104s^{3}-36s^{2}+26s+12\,.$ (161)
The explicit expression of $\mathcal{V}_{L}(F;s)$ for arbitrary $L$ is more involved and can be obtained as illustrated in the supporting Mathematica script [62].